# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the MVAR(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@
Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the LSTM(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@
Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the LSTM(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.

### v2

@@ Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@
Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@
Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the LSTM(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.